OpenStudy (deathslasher99):

State the 5 values of y corresponding to x = -2, x = -1, x = 0, x = 1, and x = 2. List each value separated by a comma (for example, -20,-10,0,10,20). y = -x + 16

OpenStudy (sshayer):

when x=-2 y=-(-2)+16=2+16=18 similarly put x=-1,0,1,2 one by one and find y

OpenStudy (deathslasher99):

So the answer is (17, 13, 9, 5, 1)